OpenStudy (anonymous):

?? [Part 1 (4 points)] Show all work in dividing 9 x squared minus 1 all over 8 x minus 4 divided by the fraction 3 x squared plus 5 x minus 2 all over 2 x squared plus x minus 1. [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (anonymous):

Just to confirm, is the problem you're solving? \[\frac{ 9x ^{2} -1 }{ 8x-4 } \div \frac{ 3x^{2}+5x-2 }{ 2x^{2} +x -1 }\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok. When you're dividing fractions, it's important to remember that dividing by a fraction is the same as multiplying by the reciprocal of that fraction. So in this example, we'll change the division sign to a multiplication sign and flip the second fraction, so we'll get this: \[\frac{ 9x^2 -1 }{ 8x -4 } \times \frac{ 2x^2 +x -1 }{ 3x^2+5x -2 }\]

OpenStudy (anonymous):

Okay.

OpenStudy (mathstudent55):

Now factor both numerators and both denominators.

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

how would you tell which one?